OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line (-4,-3) and (10, 4)?

OpenStudy (anonymous):

First need to find the slope of the line with the two points given.

OpenStudy (anonymous):

y1-y / x1-x = slopw

OpenStudy (anonymous):

4-(-3)/10-(-4) right?

OpenStudy (anonymous):

You got it. So what is the slope of that line?

OpenStudy (anonymous):

m=7/14 m=1/2

OpenStudy (anonymous):

That is correct. Now a line that is perpendicular to that line would have a slope that is the negative inverse of the slope of the original line. Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

(-1,2)?

OpenStudy (anonymous):

The slope of your line is 1/2. The negative inverse of that would be the slope of the perpendicular line. Just reverse the numbers and make it negative.

OpenStudy (anonymous):

-2/1?

OpenStudy (anonymous):

-2/1 or just -2. Good job.

OpenStudy (anonymous):

Thanks!